OpenStudy (anonymous):

a flat rectangular piece of aluminum has a perimeter of 66 inches. the length is 9 inches longer than the width. find the width. 30 inches 12 inches 21 inches 31 inches

OpenStudy (campbell_st):

let the width be w and the length w + 9 so the perimeter is P = w + w + w+9 + w +9 or P = 4w + 18 if the perimeter is 66 then 66 = 4w + 18 you now need to solve for w